Matt (Douglass) – In preparation for a proposed major shift of the circulating collection, a small shift in the East Room is being planned. As a result, some very old editions of Facts on File, Encyclopedia Britannica and World Books have been pulled and are being reviewed by the selector. This small weeding will free up one section and facilitate the shift.
Andy noted that he is re-working the proposal for the major shift, and also noted that the laptop scanner will be available to use in the withdrawal process. Perhaps a more comprehensive weeding can take place prior to the shift.
Krista (Kilmer) – Has shifted the microfilm collection to allow for more space in the cabinets.
Melanie (Math) – Continuing to work on Serials Inventory Project. Looking to free up space for more bound Periodicals. Working with Mei Ling in weeding Reserve-Math2 collection.
Megan (Physics) – The weeding of permanent Reserves and moving them to reference area has freed up space. As the collection has not been weeded in about 5 years, will begin to discuss the possibility of a weeding project.
Marty (LSM) – Nothing new to report, works continues without a hitch.
Joe (LSM) – Andy welcomed Joe Asaro. Joe is visiting all the functional group meetings to acclimate to NBACS processes and procedures.
Eddie (LSM) – Special Collections project: Busy separating items in collection into different groups depending on where they will be transferred. Will use laptop/scanner in project to transfer majority of X, X2, X-Per, and X-Folio items out of LSM. Will review all records prior to being sent to the Annex. Busy cleaning up the clutter from the weeding project, which is allowing some condensing on the shelves on the 3rd floor. Working on new stack guides. LSM CS catching up on the backlog from the weeding project.
Andy reviewed the LSM weeding project for the information of the new members and guests.
Rob (Alex)- Caught up with the end of the semester shelving, which is at a all time low in terms of returns. For example, FY98/99 - 363,021 items returned; 99/00 – 339,793; 00/01 – 303,069; 01/02- 312,675; 02/03 – 292,350;03/04 – 276,781. The mold in the 3rd floor ceiling was corrected over the winter break; staff never lost access to the 3rd floor while the work was on going. The crack has been repaired but the problem in the Mechanical Room has not, so anticipate more leaks in future because of this. Reference re-label project is completed. Working on updating floor maps and stack guides.
Penny – Noted that Douglass year-end returns were also light. Each staff was scheduled a hour of shelving duties, thus clearing up the returns before the end of the first week back from break. The few remaining trucks were completed by Dglss students working over break.
Stacy (Annex) - Temperature and humidity measurements have leveled off. Humidity controls have been shut off and the humidity measurements remain a constant 30%. Stacy has created charts and graphs depicting discrepancies and will continue to monitor due to continued power fluxes that throw the humidity levels off. As a result, Stacy is concerned about mold. She questioned the group regarding re-con projects in terms of missing books. Please notify her when working on these projects so Annex staff can search, as she believes some of these items maybe at the Annex. Continues to receive transfers from LSM weeding project, which is going well. Working on backlog of Bound Periodicals.
Shirley (Alex) – Working on In-Process materials. The Missing books lists are getting smaller, possibly due in part to the In-transit lists. These lists are wonderful as communication between units is good and helps in clearing the lists.
The final draft of the proposal has been forwarded to Judy and Roger and has received preliminary approval. Once final approval is received, a condensed version will be placed on the NBL page for group members to use at their units. The document contains condensed instructions for the "LC Call Number" and "Shelved by Title" Range Guide Templates as well as both templates. Melanie offered to test the document to see if it can become a merge file.
Of the original 112 items charged to the list of internal users, 27 items remain checked out. Andy distributed an excel spreadsheet handout listing the 27 items and requested that the group search their units one last time. Andy will send a copy via e-mail attachment so members can sort as they see fit (e.g. call # order). If an item is located, please discharge it until it is in-transit to its owning library and mark the spreadsheet with an "F". If not found, mark the spreadsheet with a "NF". Please complete your searches and get the spreadsheet back to Andy by Thursday, January 27th. This type of project, where the group works with reports to correct a problem, gives the group a flavor of what Andy hopes the group can work on and accomplish in the upcoming year.
Andy distributed a DRAFT of the report for group review. Discussion ensued regarding future goals most particularly related to the group working on projects driven by reports. Andy requested input on any necessary additions of activities from 8/04 to date, that he may have missed in the DRAFT. After any additional information has been added he will submit the report. The additional goals being considered are: Tracking staff time spent in searching; group attendance at applicable workshops; NBL standardized floor plans/maps; site tours.
Wednesday, February 9, 2005